Joe348:

A helicopter flying 3,590 feet above ground spots the top of a 150-foot tall tower at an angle of depression of 77° . How far must the helicopter fly to be directly over the tower?

aidenj:

what do u have so far

Joe348:

um not much

Joe348:

Jus that ^

Joe348:

why

aidenj:

what do u think we need to do

Joe348:

idek

aidenj:

can u try

Joe348:

I looked it up and it said like subtractions first

aidenj:

so what do we do from there what do we subtract

Joe348:

um it says 3590-150

aidenj:

ok what is that

Joe348:

3400

Joe348:

x = 3440 / tan(77)

aidenj:

ok do that

Joe348:

794.18

Joe348:

but idek if thats right

snowflake0531:

Yea it wouldbe tan(77) = (3590-150)/x So yea this is correct

@joe348 wrote:
x = 3440 / tan(77)

Joe348:

Okay thanjs snow

Joe348:

thanks*

snowflake0531:

oc oc

aaronlloyd:

aaronlloyd: